OpenStudy (anonymous):

How to find out if this an identity? sec x + csc x = 1 Help Please, I have no clue.

OpenStudy (anonymous):

you can show that it is not (because it certainly is not) by picking a value for \(x\) and see that what you get does not add up to one

OpenStudy (anonymous):

pick something easy, like \(\frac{\pi}{4}\)

OpenStudy (anonymous):

them compute \[\sec(\frac{\pi}{4})+\sin(\frac{\pi}{4})\] and see that you do not get 1